# Vertex-minimal hyperbolic origami 2-torus

### v1

@@ 3.1 Robust Flatness
Next, we explain how we otain the cone angle bounds from $\alpha_{i,j}$ and $\beta_{i,j}$. Consider two functions $\phi_{+}$ and $\phi_{-}$ given by $\phi_{+}(x)=\arccos(\sqrt{x})$ and $\phi_{-}(x)=\arccos(-\sqrt{x})$. Denote by $\theta_{i,j}$ the hyperbolic angle at $\hat{X}_{i}$ of the triangle $(\hat{X}_{i},X_{n_{j}},\hat{X}_{n_{j+1}})$, and denote by $\tau_{i,j}$ the Euclidean angle between the two vectors $Y_{i,n_{j}}$ and $Y_{i,n_{j+1}}$. Since $\langle X_{n_{j}}-X_{i},X_{n_{j+1}}-X_{i}\rangle_{X_{i}}$ and $\langle Y_{i,n_{j}},Y_{i,n_{j+1}}\rangle$ always share the same sign, we know that
@@ 3.2 Robust Embeddedness
The next step is to show that if we slightly purturb the $z$-coordinates of $\hat{S}$, the resulting 2-torus is still embedded. The key observation (and the main reason we use the Beltrami–Klein model of the hyperbolic plane) is that the underlying set of the hyperbolic geodesic triangle with vertices $V_{1},V_{2},V_{3}\in\mathbb{H}^{3}$ is precisely the Euclidean triangle with the same vertices. Therefore, in this section we “forget” the hyperbolic structure and instead treat $\hat{S}$ as a map $\hat{S}:(\Sigma_{2},\mathcal{T})\rightarrow B_{1}(0)$ whose image is the unit Euclidean ball. We begin with the following definition, which is not “natural” in the hyperbolic setting but is more appropriate for Euclidean geometry.
@@
A map $S:(\Sigma_{2},\mathcal{T})\rightarrow B_{1}(0)$ is $\lambda$-robustly embedded if for all $S^{\prime}$ obtained from moving each of the $z$-coordinate of the vertices of $S$ by at most $\lambda$ is also embedded.
@@
We know that every $X\in T$ can be expressed as $X=t_{1}V_{1}+t_{2}V_{2}+t_{3}V_{3}$ where $t_{i}\geq 0$ and $t_{1}+t_{2}+t_{3}=1$. Linearity of the dot product then implies the minimum of $\langle X,\hat{N}\rangle$ must be achieved at one of the vertices of $T$, so Equation (10) implies (9). Now, if $T^{\prime}_{1}$ and $T^{\prime}_{2}$ are obtained from purturbing the $z$-coordinates of the vertices of $T_{1}$ and $T_{2}$ by at most $\delta$, then the dot product $\langle V_{i},\hat{N}\rangle$ is purturbed by at most $\delta C$. It follows that
@@
Suppose $T^{\prime}_{1}$ and $T^{\prime}_{2}$ are obtained from purturbing $(U,V_{1},V_{2},W_{1},W_{2})$ by at most $\delta$. The dot products with $\hat{N}$ is purturbed by at most $\delta C$. It follows that
@@
$\langle U^{\prime},\hat{N}\rangle-\langle X^{\prime},\hat{N}\rangle=s_{1}\langle U^{\prime},-W^{\prime}_{1},\hat{N}\rangle+s_{2}\langle U^{\prime}-W^{\prime}_{2},\hat{N}\rangle$.
@@
$t_{1}\langle V^{\prime}_{1}-U^{\prime},\hat{N}\rangle+t_{2}\langle V^{\prime}_{2}-U^{\prime},\hat{N}\rangle+s_{1}\langle U^{\prime},-W^{\prime}_{1},\hat{N}\rangle+s_{2}\langle U^{\prime}-W^{\prime}_{2},\hat{N}\rangle=0,$
@@
Among the 276 pairs of triangles from $F$, there are 82 pairs that share no vertices in common, and 158 of which share one vertex in common. We provide an algorithm to generate $\hat{N}$ for each of these pairs where the conditions in Lemma 3.6 and 3.7 holds. Consider a map $\rho:\mathbb{Z}_{>0}\rightarrow\mathbb{Z}^{3}$ as follows:
@@
Lemma 3.5 then implies the $10^{32}$-diluted surface $\hat{S}^{\prime}$ is $10^{25}$-robustly enbedded, so the original surface $\hat{S}$ is $10^{7}$-robustly embedded, as desired. ∎
@@ 4 Finding Origami 2-Tori
Let $B_{r}(p)\subset\mathbb{R}^{n}$ denote the open ball of radius $r$ centered at $p$. If $G:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is $\lambda$-expansive on $B_{r}(p)$ for some $\lambda>0$, then $B_{\lambda r}(G(p))\subset G(B_{r}(p))$.
@@ 4.1 The Jacobian
Our first goal is to analytically computing the Jacobian matrix of the map $\Theta$. Recall from Equation (4) that $\Theta$ is given by
@@
where $\theta_{ijk}(z)$ denote the angle of the triangle $(i,j,k)$ at vertex $i$.
@@
$\left|\frac{\partial\Theta_{i}}{\partial x_{j}}(z)-\frac{\partial\Theta_{i}}{\partial x_{j}}(\hat{z})\right|\leq 10\cdot r\cdot\sup_{k\in[10],w\in B}\left|\frac{\partial^{2}\Theta_{i}}{\partial z_{j}\partial z_{k}}(w)\right|<10^{-3}.$
@@
We explain how we find that the singular values of $M$ are greater than $1.5$. Since the entries of $M^{t}M$ are obtained from taking sums and products of entries of $M$, we may compute them precisely using floating-point arithmetic. This further implies that we may compute precisely the coefficients of the characteristic polynomial $\chi$. We find that the characteristic polynomial has 10 distinct roots $x_{i}$ by finding $a_{i},b_{i}\in\mathbb{R}$ such that $a_{i}<b_{i}$ and $\chi(a_{i})\chi(b_{i})<0$. The intermediate value theorem then guarantees the existence of the root $x_{i}\in(a_{i},b_{i})$ for each of the ten pairs $(a_{i},b_{i})$. We omit the tedius computation.
@@ 4.2 Bounds on Angles and Hyperbolic Distances
The second inequaity follows from a routine application of the Cauchy-Schwarz inequality and monotonicity of $1/(1-x^{2})^{2}$ on $[0,r]$:
@@
Denote by $\hat{V}=\hat{Y}-\hat{X}$, where $\hat{V}$ is an edge connecting $\hat{X}$ to $\hat{Y}$ in $\hat{S}$ that corresponds to $V$. We compute numerically using Equation (5) that $\|V\|\in[0.509,1.561]$. Then, since $\hat{S}\subset B_{0.79}(0)$, we must have $S\subset B_{0.8}(0)$ for all $S\in\mathcal{B}$, so applying Lemma 4.7 with $r=0.8$ gives us the desired bounds. ∎
@@
Using the coordinates of $\hat{S}$, we compute that $l(\hat{e})\in[0.63,2.08]$. We explain how we establish the bound. Given two distinct points $X,Y\in\mathbb{K}^{3}$, the quadratic polynomial $\|(1-t)X+tY\|^{2}=1$ has two distinct real roots $t_{1}<0$ and $t_{2}>1$. Let $A=(1-t_{1})X+t_{1}Y$ and $B=(1-t_{2})X+t_{2}Y$. The hyperbolic distance $d_{g}(X,Y)$ is given by
@@ 4.3 Bounds on the Second-Order Partials
  • $|z_{i}|\leq 0.8$ ($S\subset B_{0,8}(0)$).
We remark that the Cauchy-Schwarz inequality further implies $u\leq 1.6^{2}$.
@@
$|\partial_{l}\theta|=\frac{\partial_{l}(vw)\cos\theta-\partial_{l}u}{vw\sin\theta}\leq\frac{2\cdot 1.6\cdot 10^{3}+10^{3}}{0.5^{2}\cdot 0.34}\leq 3.2\cdot 10^{6}\leq 10^{7}.$
@@
To ease notations we write

### v2

@@ 3.1 Robust Flatness
Next, we explain how we obtain the cone angle bounds from $\alpha_{i,j}$ and $\beta_{i,j}$. Consider two functions $\phi_{+}$ and $\phi_{-}$ given by $\phi_{+}(x)=\arccos(\sqrt{x})$ and $\phi_{-}(x)=\arccos(-\sqrt{x})$. Denote by $\theta_{i,j}$ the hyperbolic angle at $\hat{X}_{i}$ of the triangle $(\hat{X}_{i},X_{n_{j}},\hat{X}_{n_{j+1}})$, and denote by $\tau_{i,j}$ the Euclidean angle between the two vectors $Y_{i,n_{j}}$ and $Y_{i,n_{j+1}}$. Since $\langle X_{n_{j}}-X_{i},X_{n_{j+1}}-X_{i}\rangle_{X_{i}}$ and $\langle Y_{i,n_{j}},Y_{i,n_{j+1}}\rangle$ always share the same sign, we know that
@@ 3.2 Robust Embeddedness
The next step is to show that if we slightly perturb the $z$-coordinates of $\hat{S}$, the resulting 2-torus is still embedded. The key observation (and the main reason we use the Beltrami–Klein model of $\mathbb{H}^{3}$) is that the underlying set of the hyperbolic geodesic triangle with vertices $V_{1},V_{2},V_{3}\in\mathbb{H}^{3}$ is precisely the Euclidean triangle with the same vertices. Therefore, in this section we “forget” the hyperbolic structure and instead treat $\hat{S}$ as a map $\hat{S}:(\Sigma_{2},\mathcal{T})\rightarrow B_{1}(0)$ whose image is the unit Euclidean ball. We begin with the following definition, which is not “natural” in the hyperbolic setting but is more appropriate for Euclidean geometry.
@@
A map $S:(\Sigma_{2},\mathcal{T})\rightarrow B_{1}(0)$ is $\lambda$-robustly embedded if for all $S^{\prime}$ obtained from moving each of the $z$-coordinates of the vertices of $S$ by at most $\lambda$ is also embedded.
@@
We know that every $X\in T$ can be expressed as $X=t_{1}V_{1}+t_{2}V_{2}+t_{3}V_{3}$ where $t_{i}\geq 0$ and $t_{1}+t_{2}+t_{3}=1$. Linearity of the dot product then implies the minimum of $\langle X,\hat{N}\rangle$ must be achieved at one of the vertices of $T$, so Equation (11) implies (10). Now, if $T^{\prime}_{1}$ and $T^{\prime}_{2}$ are obtained from purturbing the $z$-coordinates of the vertices of $T_{1}$ and $T_{2}$ by at most $\delta$, then the dot product $\langle V_{i},\hat{N}\rangle$ is perturbed by at most $\delta C$. It follows that
@@
Suppose $T^{\prime}_{1}$ and $T^{\prime}_{2}$ are obtained from purturbing $(U,V_{1},V_{2},W_{1},W_{2})$ by at most $\delta$. The dot products with $\hat{N}$ is perturbed by at most $\delta C$. It follows that
@@
$\langle U^{\prime},\hat{N}\rangle-\langle X^{\prime},\hat{N}\rangle=s_{1}\langle U^{\prime}-W^{\prime}_{1},\hat{N}\rangle+s_{2}\langle U^{\prime}-W^{\prime}_{2},\hat{N}\rangle$.
@@
$t_{1}\langle V^{\prime}_{1}-U^{\prime},\hat{N}\rangle+t_{2}\langle V^{\prime}_{2}-U^{\prime},\hat{N}\rangle+s_{1}\langle U^{\prime}-W^{\prime}_{1},\hat{N}\rangle+s_{2}\langle U^{\prime}-W^{\prime}_{2},\hat{N}\rangle=0,$
@@
Among the 276 pairs of triangles from $F$, there are 82 pairs that share no vertices in common, and 158 of which share one vertex in common. We provide an algorithm to generate $\hat{N}$ for each of these pairs where the conditions in Lemma 3.6 and 3.7 hold. Consider a map $\rho:\mathbb{Z}_{>0}\rightarrow\mathbb{Z}^{3}$ as follows:
@@
Lemma 3.5 then implies the $10^{32}$-dilated surface $\hat{S}^{\prime}$ is $10^{25}$-robustly embedded, so the original surface $\hat{S}$ is $10^{7}$-robustly embedded, as desired. ∎
@@ 4 Finding Origami 2-Tori
Let $B_{r}(p)\subset\mathbb{R}^{n}$ denote the open ball of radius $r$ centered at $p$. If $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is $\lambda$-expansive on $B_{r}(p)$ for some $\lambda>0$, then $B_{\lambda r}(f(p))\subset f(B_{r}(p))$.
@@ 4.1 The Jacobian
Our first goal is to analytically compute the Jacobian matrix of the map $\Theta$. Recall from Equation (4) that $\Theta$ is given by
@@
where $\theta_{ijk}(z)$ denotes the angle of the triangle $(i,j,k)$ at vertex $i$.
@@
$\left|\frac{\partial\Theta_{i}}{\partial z_{j}}(z)-\frac{\partial\Theta_{i}}{\partial z_{j}}(\hat{z})\right|\leq 10\cdot r\cdot\sup_{k\in[10],w\in B}\left|\frac{\partial^{2}\Theta_{i}}{\partial z_{j}\partial z_{k}}(w)\right|<10^{-3}.$
@@
We explain how we find that the singular values of $M$ are greater than $1.5$. Since the entries of $M^{t}M$ are obtained from taking sums and products of entries of $M$, we may compute them precisely using floating-point arithmetic. This further implies that we may compute precisely the coefficients of the characteristic polynomial $\chi$. We find that the characteristic polynomial has 10 distinct roots $x_{i}$ by finding $a_{i},b_{i}\in\mathbb{R}$ such that $a_{i}<b_{i}$ and $\chi(a_{i})\chi(b_{i})<0$. The intermediate value theorem then guarantees the existence of the root $x_{i}\in(a_{i},b_{i})$ for each of the ten pairs $(a_{i},b_{i})$. We omit the tedious computation.
@@ 4.2 Bounds on Angles and Hyperbolic Distances
The second inequality follows from a routine application of the Cauchy-Schwarz inequality and monotonicity of $1/(1-x^{2})^{2}$ on $[0,r]$:
@@
Denote by $\hat{V}=\hat{Y}-\hat{X}$, where $\hat{V}$ is an edge connecting $\hat{X}$ to $\hat{Y}$ in $\hat{S}$ that corresponds to $V$. We compute numerically using Equation (6) that $\|\hat{V}\|\in[0.509,1.561]$. Then, since $\hat{S}\subset B_{0.79}(0)$, we must have $S\subset B_{0.8}(0)$ for all $S\in\mathcal{B}$, so applying Lemma 4.7 with $r=0.8$ gives us the desired bounds. ∎
@@
Using the coordinates of $\hat{S}$, we compute that $l(\hat{e})\in[0.63,2.08]$. We explain how we establish the bound. Given two distinct points $X,Y\in\mathbb{H}^{3}$, the quadratic polynomial $\|(1-t)X+tY\|^{2}=1$ has two distinct real roots $t_{1}<0$ and $t_{2}>1$. Let $A=(1-t_{1})X+t_{1}Y$ and $B=(1-t_{2})X+t_{2}Y$. The hyperbolic distance $d_{g}(X,Y)$ is given by
@@ 4.3 Bounds on the Second-Order Partials
  • $|z_{i}|\leq 0.8$ ($S\subset B_{0.8}(0)$).
We remark that the Cauchy-Schwarz inequality further implies $|u|\leq 1.6^{2}$.
@@
$|\partial_{l}\theta|=\frac{\partial_{l}(vw)\cos\theta-\partial_{l}u}{vw\sin\theta}\leq\frac{2\cdot 1.6\cdot 10^{3}+10^{3}}{0.5^{2}\cdot 0.24}\leq 3.2\cdot 10^{6}\leq 10^{7}.$
@@
To ease notation we write